OpenStudy (anonymous):

Find the critical numbers of the function on the interval 0 ≤ θ < 2π. f(θ) = 2cos(θ) + (sin(θ))^2 θ = (smaller value) θ = (larger value)

myininaya (myininaya):

Find f'.

myininaya (myininaya):

Have you done that yet?

myininaya (myininaya):

I will give you this: \[(\cos(\theta))'=-\sin(\theta)\] \[(\sin(\theta))'=\cos(\theta)\] \[([g(\theta)]^2)'=2[g(\theta)]^{2-1}=2g(\theta) \]

OpenStudy (anonymous):

Yes but im having trouble when i make the derivative equal to 0

myininaya (myininaya):

So you got: \[f'(\theta)=-2\sin(\theta)+2\sin(\theta)(\cos(\theta)) ?\]

myininaya (myininaya):

Are you still there @dainel40 ?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

yes I got that

OpenStudy (anonymous):

@myininaya still there?

myininaya (myininaya):

Ok set equal to 0 Look and see that both terms have the common factor sin(theta)

myininaya (myininaya):

then factor the expression using that.

OpenStudy (anonymous):

factor out a \(2\sin(\theta)\)

OpenStudy (anonymous):

hello! i didn't see you

myininaya (myininaya):

I was away for a sec.

OpenStudy (anonymous):

im getting f'(0) = pi not sure if its right

myininaya (myininaya):

\[f'(\theta)=0 \] when theta=? Did you factor the expression using what satellite or I said? Either way will work but satellite's looks prettier.

myininaya (myininaya):

You should get two values.

OpenStudy (anonymous):

yes, im getting pi

OpenStudy (anonymous):

umm

myininaya (myininaya):

\[2 \sin(\theta)(-1+\cos(\theta))=0\] ? Setting both equal to 0? Remember you are looking at the interval [0,2pi)

OpenStudy (anonymous):

ok im getting pi, 2pi now

myininaya (myininaya):

I define a critical number c as a number that satisfies f'(c)=0 or f'(c) dne and also c exists as an element of the domain of f.

myininaya (myininaya):

Does 2pi exist in the domain of f?

OpenStudy (anonymous):

no because it is not included

myininaya (myininaya):

So our only critical numbers are?

OpenStudy (anonymous):

pi

myininaya (myininaya):

and ?

myininaya (myininaya):

there is another number you can plug in that will give you 0

myininaya (myininaya):

what is it

OpenStudy (anonymous):

0?

myininaya (myininaya):

yep.

myininaya (myininaya):

so are critical numbers are 0 and pi

OpenStudy (anonymous):

yay that was correct!!!

OpenStudy (anonymous):

thanks alot!!!

myininaya (myininaya):

lol

myininaya (myininaya):

np